# Sharp non-uniqueness for the Navier–Stokes equations in scaling critical spaces

### v1

@@ 3. Highly oscillating functions in Besov spaces
$F_{\lambda,L}^{({\rm p})}:=\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}g_{\lambda,\ell,k},$ (3.31)
$\begin{aligned} F_{\lambda,L}^{({\rm r})}:={}&\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{\begin{subarray}{c}0\leqslant k^{\prime}\leqslant K\\ k^{\prime}\neq k\end{subarray}}\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}\\ &+\sum_{k^{\prime}=0}^{K}\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\geqslant 3}P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}.\end{aligned}$ (3.32)
@@
${\left\|F_{\lambda,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant\sum_{k=1}^{K}\left\{\sum_{j^{\prime}\in\mathbb{Z}}\left(2^{sj}{\left\|\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.33)
$=\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell,k}|\leqslant 3}\left(2^{sj}{\left\|\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.34)
$\leqslant C\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell,k}|\leqslant 3}\left(2^{sj}{\left\|g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.35)
$=C\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell,k}|\leqslant 3}\left(2^{sj}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}{\left\|f_{k}\right\|}_{L^{p}}$ (3.36)
$\leqslant C\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}.$ (3.37)
@@
${\left\|F_{\lambda,L}^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{s}}$ $\leqslant\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{\begin{subarray}{c}0\leqslant k^{\prime}\leqslant K\\ k^{\prime}\neq k\end{subarray}}\sum_{|j-j_{\ell,k}|\leqslant 2}\sum_{|j^{\prime}-j|\leqslant 1}2^{sj^{\prime}}{\left\|P_{k^{\prime}}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.38)
$\qquad+\sum_{k=1}^{K}\sum_{k^{\prime}=0}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\geqslant 3}\sum_{|j^{\prime}-j|\leqslant 1}2^{sj^{\prime}}{\left\|P_{k^{\prime}}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.39)
$\leqslant C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(k^{\prime},j)\in R_{\ell,k}}2^{sj}{\left\|P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}\right\|}_{L^{p}}.$ (3.40)
@@
${\left\|P_{m}\Delta_{j}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ $\leqslant|b_{\ell}|\sum_{j^{\prime}\geqslant\max\{j,j_{\ell,k}\}+j_{*}}{\left\|\Delta_{j^{\prime}}\chi_{m,j,\ell,k}*f_{k}\right\|}_{L^{p}}$ (3.41)
$\leqslant C|b_{\ell}|\sum_{j^{\prime}\geqslant\max\{j,j_{\ell,k}\}+j_{*}}2^{-Mj^{\prime}}{\left\|\chi_{m,j,\ell,k}*\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.42)
$=C|b_{\ell}|\left(\max\{2^{j},\lambda^{\ell}|a_{k}|\}\right)^{-M}{\left\|\chi_{m,j,\ell,k}*\nabla^{M}f_{k}\right\|}_{L^{p}},$ (3.43)
where we have put $j_{*}=\log_{2}(2c_{*})$. Then, we obtain
${\left\|F_{\lambda,L}^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{s}}\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(k^{\prime},j)\in R_{\ell,k}}|b_{\ell}|2^{sj}\left(\max\{2^{j},\lambda^{\ell}|a_{k}|\}\right)^{-M}{\left\|\chi_{k^{\prime},j,\ell,k}*\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.44)
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{k^{\prime}=0}^{K}\sum_{j\geqslant j_{\ell,k}+1}|b_{\ell}|2^{(s-M)j}{\left\|\chi_{k^{\prime},j,\ell,k}\right\|}_{L^{1}}{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.45)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{k^{\prime}=0}^{K}\sum_{j\leqslant j_{\ell,k}}|b_{\ell}|2^{sj}(\lambda^{\ell}|a_{k}|)^{-M}{\left\|\chi_{k^{\prime},j,\ell,k}\right\|}_{L^{p}}{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}$ (3.46)
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j\geqslant j_{\ell,k}+1}|b_{\ell}|2^{(s-M)j}{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.47)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j\leqslant j_{\ell,k}}|b_{\ell}|2^{\{s+n(1-\frac{1}{p})\}j}(\lambda^{\ell}|a_{k}|)^{-M}{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}$ (3.48)
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.49)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}.$ (3.50)
@@
${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant c\max_{1\leqslant k\leqslant K}{\left\|P_{k}F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}.$ (3.51)
Fix a $k\in\{1,\dots,K\}$ and decompose $P_{k}F_{\lambda,L}$ as
$P_{k}F_{\lambda,L}=F^{({\rm p})}_{\lambda,k,L}+F^{({\rm r})}_{\lambda,k,L}$; (3.52)
$\quad F^{({\rm p})}_{\lambda,k,L}:=\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}g_{\lambda,\ell,k},$ (3.53)
$\quad F^{({\rm r})}_{\lambda,k,L}:=\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\geqslant 3}P_{k}\Delta_{j}g_{\lambda,\ell,k}+\sum_{\begin{subarray}{c}1\leqslant k^{\prime}\leqslant K\\ k^{\prime}\neq k\end{subarray}}\sum_{\ell=1}^{L}P_{k}g_{\lambda,\ell,k^{\prime}}.$ (3.54)
For the lower bound estimate of the principal part $F^{({\rm p})}_{\lambda,k,L}$, we see that
${\left\|F_{\lambda,k,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}$ $=\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell,k}|\leqslant 3}\left(2^{sj^{\prime}}{\left\|\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.55)
$\geqslant\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\left(2^{sj_{\ell,k}}{\left\|\sum_{|j-j_{\ell,k}|\leqslant 2}P_{k}\Delta_{j}\Delta_{j_{\ell,k}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.56)
$=\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\left(2^{sj_{\ell,k}}{\left\|P_{k}\Delta_{j_{\ell,k}}g_{\lambda,\ell,k}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}.$ (3.57)
@@
$P_{k}\Delta_{j_{\ell,k}}g_{\lambda,\ell,k}(x)$ $=b_{\ell}\int_{\mathbb{R}^{n}}\chi_{k,j_{\ell,k}}(y)f_{k}(x-y)e^{i\lambda^{\ell}a_{k}\cdot(x-y)}dy$ (3.58)
$=b_{\ell}e^{i\lambda^{\ell}a_{k}\cdot x}f_{k}(x)\int_{\mathbb{R}^{n}}\chi_{k,j_{\ell,k}}(y)e^{-i\lambda^{\ell}a_{k}\cdot y}dy$ (3.59)
$\qquad+b_{\ell}e^{i\lambda^{\ell}a_{k}\cdot x}\int_{\mathbb{R}^{n}}(f_{k}(x-y)-f_{k}(x))\chi_{k,j_{\ell,k}}(y)e^{-i\lambda^{\ell}a_{k}\cdot y}dy$ (3.60)
$=:b_{\ell}e^{i\lambda^{\ell}a_{k}\cdot x}f_{k}(x)+R_{\lambda,\ell,k}(x),$ (3.61)
where we have used $\widehat{\chi_{k,j_{\ell,k}}}(\lambda^{\ell}a_{k})=1$. The remainder term $R_{\lambda,\ell,k}$ is bounded as
${\left\|R_{\lambda,\ell,k}\right\|}_{L^{p}}$ $\leqslant|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}\int_{\mathbb{R}^{n}}|y||\chi_{k,j_{\ell,k}}(y)|dy$ (3.62)
$\leqslant C|b_{\ell}|2^{-j_{\ell,k}}{\left\|\nabla f_{k}\right\|}_{L^{p}}$ (3.63)
$\leqslant C|b_{\ell}|(\lambda^{\ell}|a_{k}|)^{-1}{\left\|\nabla f_{k}\right\|}_{L^{p}}.$ (3.64)
@@
${\left\|P_{k}\Delta_{j_{\ell,k}}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ $\geqslant|b_{\ell}|{\left\|f_{k}\right\|}_{L^{p}}-C(\lambda^{\ell}|a_{k}|)^{-1}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}},$ (3.65)
which implies
${\left\|F_{\lambda,k,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $c\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\max_{1\leqslant k\leqslant K}|a_{k}|^{s}{\left\|f_{k}\right\|}_{L^{p}}$ (3.66)
$-C\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s-1}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}.$ (3.67)
On the estimate of $F^{({\rm r})}_{\lambda,k,L}$, the similar strategy as for the estimate of $F^{({\rm r})}_{\lambda,k,L}$ yields
${\left\|F^{({\rm r})}_{\lambda,k,L}\right\|}_{\dot{B}_{p,q}^{s}}\leqslant{}$ $\sum_{\ell=1}^{L}\sum_{|j-j_{\ell,k}|\geqslant 3}\sum_{|j^{\prime}-j|\leqslant 1}2^{sj^{\prime}}{\left\|P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.68)
$+\sum_{\begin{subarray}{c}1\leqslant k^{\prime}\leqslant K\\ k^{\prime}\neq k\end{subarray}}\sum_{\ell=1}^{L}\sum_{j\in\mathbb{Z}}2^{sj}{\left\|P_{k}\Delta_{j}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.69)
$\leqslant{}$ $C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}\sum_{(k^{\prime},j)\in R_{\ell,k}}2^{sj}{\left\|P_{k^{\prime}}\Delta_{j}g_{\lambda,\ell,k}\right\|}_{L^{p}}$ (3.70)
$\leqslant{}$ $C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.71)
$+C\max_{1\leqslant k\leqslant K}\sum_{\ell=1}^{L}(\lambda^{\ell}|a_{k}|)^{s+n(1-\frac{1}{p})-M}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}.$ (3.72)
Gathering (3.67) and (3.72), we complete the proof. ∎

### v2

@@ 3. Highly oscillating functions in Besov spaces
$F_{\lambda,L}^{({\rm p})}:=\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{|j-j_{\ell}|\leqslant 2}P_{k}\Delta_{j}g_{\lambda,k,\ell},$ (3.31)
$\begin{aligned} F_{\lambda,L}^{({\rm r})}:={}&\sum_{m=0}^{K}\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(m,j)\in R_{k,\ell}}P_{m}\Delta_{j}g_{\lambda,k,\ell}.\end{aligned}$ (3.32)
@@
${\left\|F_{\lambda,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant\sum_{k=1}^{K}\left\{\sum_{j^{\prime}\in\mathbb{Z}}\left(2^{sj^{\prime}}{\left\|\sum_{\ell=1}^{L}\sum_{|j-j_{\ell}|\leqslant 1}P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,k,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.33)
$=\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell}|\leqslant 2}\left(2^{sj}{\left\|\sum_{|j-j_{\ell}|\leqslant 1}P_{k}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,k,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.34)
$\leqslant C\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell}|\leqslant 2}\left(2^{sj}{\left\|g_{\lambda,k,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.35)
$=C\sum_{k=1}^{K}\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell}|\leqslant 2}\left(2^{sj}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}{\left\|f_{k}\right\|}_{L^{p}}$ (3.36)
$\leqslant C\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}\sum_{k=1}^{K}{\left\|f_{k}\right\|}_{L^{p}}.$ (3.37)
@@
${\left\|F_{\lambda,L}^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{s}}$ $\leqslant\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(m,j)\in R_{k,\ell}}2^{sj^{\prime}}{\left\|P_{k^{\prime}}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,k,\ell}\right\|}_{L^{p}}$ (3.38)
$\leqslant C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(k^{\prime},j)\in R_{k,\ell}}2^{sj}{\left\|P_{k^{\prime}}\Delta_{j}g_{\lambda,k,\ell}\right\|}_{L^{p}}.$ (3.39)
@@
${\left\|P_{m}\Delta_{j}g_{\lambda,k,\ell}\right\|}_{L^{p}}$ $\leqslant|b_{\ell}|\sum_{j^{\prime}\geqslant\max\{j,j_{\ell}\}+j_{*}}{\left\|\Delta_{j^{\prime}}\chi_{m,j,k,\ell}*f_{k}\right\|}_{L^{p}}$ (3.40)
$\leqslant C|b_{\ell}|\sum_{j^{\prime}\geqslant\max\{j,j_{\ell}\}+j_{*}}2^{-Mj^{\prime}}{\left\|\chi_{m,j,k,\ell}*\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.41)
$=C|b_{\ell}|\left(\max\{2^{j},\lambda^{\ell}\}\right)^{-M}{\left\|\chi_{m,j,k,\ell}*\nabla^{M}f_{k}\right\|}_{L^{p}},$ (3.42)
where we have put $j_{*}=\log_{2}(2c_{*})$. Thus, we obtain
${\left\|F_{\lambda,L}^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{s}}\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{(m,j)\in R_{k,\ell}}|b_{\ell}|2^{sj}\left(\max\{2^{j},\lambda^{\ell}\}\right)^{-M}{\left\|\chi_{m,j,k,\ell}*\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.43)
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{m=0}^{K}\sum_{j\geqslant j_{\ell}+1}|b_{\ell}|2^{(s-M)j}{\left\|\chi_{m,j,k,\ell}\right\|}_{L^{1}}{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.44)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{m=0}^{K}\sum_{j\leqslant j_{\ell}}|b_{\ell}|2^{sj}(\lambda^{-M\ell}{\left\|\chi_{m,j,k,\ell}\right\|}_{L^{p}}{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}$ (3.45)
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j\geqslant j_{\ell}+1}|b_{\ell}|2^{(s-M)j}{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.46)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j\leqslant j_{\ell}}|b_{\ell}|2^{\{s+n(1-\frac{1}{p})\}j}\lambda^{-M\ell}{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}$ (3.47)
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.48)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s+n(1-\frac{1}{p})-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{1}}.$ (3.49)
@@
${\left\|F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant c\max_{1\leqslant m\leqslant K}{\left\|P_{m}F_{\lambda,L}\right\|}_{\dot{B}_{p,q}^{s}}.$ (3.50)
Fix a $m\in\{1,\dots,K\}$ and decompose $P_{m}F_{\lambda,L}$ as
$P_{m}F_{\lambda,L}=F^{({\rm p})}_{\lambda,m,L}+F^{({\rm r})}_{\lambda,m,L}$; (3.51)
$\quad F^{({\rm p})}_{\lambda,m,L}:=\sum_{\ell=1}^{L}\sum_{|j-j_{\ell}|\leqslant 1}P_{m}\Delta_{j}g_{\lambda,m,\ell},$ (3.52)
$\quad F^{({\rm r})}_{\lambda,m,L}:=\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j:(m,j)\in R_{k,\ell}}P_{m}\Delta_{j}g_{\lambda,k,\ell}.$ (3.53)
For the lower bound estimate of the principal part $F^{({\rm p})}_{\lambda,m,L}$, we see that
${\left\|F_{\lambda,m,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}$ $=\left\{\sum_{j^{\prime}\in\mathbb{Z}}\left(2^{sj^{\prime}}{\left\|\sum_{\ell=1}^{L}\sum_{|j-j_{\ell}|\leqslant 1}P_{m}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,m,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.54)
$=\left\{\sum_{\ell=1}^{L}\sum_{|j^{\prime}-j_{\ell}|\leqslant 2}\left(2^{sj^{\prime}}{\left\|\sum_{|j-j_{\ell}|\leqslant 1}P_{m}\Delta_{j}\Delta_{j^{\prime}}g_{\lambda,m,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.55)
$\geqslant\left\{\sum_{\ell=1}^{L}\left(2^{sj_{\ell}}{\left\|\sum_{|j-j_{\ell}|\leqslant 1}P_{m}\Delta_{j}\Delta_{j_{\ell}}g_{\lambda,m,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}$ (3.56)
$=\left\{\sum_{\ell=1}^{L}\left(2^{sj_{\ell}}{\left\|P_{m}\Delta_{j_{\ell}}g_{\lambda,m,\ell}\right\|}_{L^{p}}\right)^{q}\right\}^{\frac{1}{q}}.$ (3.57)
@@
$P_{m}\Delta_{j_{\ell}}g_{\lambda,m,\ell}(x)$ $=b_{\ell}\int_{\mathbb{R}^{n}}\chi_{m,j_{\ell}}(y)f_{m}(x-y)e^{i\lambda^{\ell}a_{m}\cdot(x-y)}dy$ (3.58)
$=b_{\ell}e^{i\lambda^{\ell}a_{m}\cdot x}f_{m}(x)\int_{\mathbb{R}^{n}}\chi_{m,j_{\ell}}(y)e^{-i\lambda^{\ell}a_{m}\cdot y}dy$ (3.59)
$\qquad+b_{\ell}e^{i\lambda^{\ell}a_{k}\cdot x}\int_{\mathbb{R}^{n}}(f_{m}(x-y)-f_{m}(x))\chi_{m,j_{\ell}}(y)e^{-i\lambda^{\ell}a_{m}\cdot y}dy$ (3.60)
$=:b_{\ell}e^{i\lambda^{\ell}a_{m}\cdot x}f_{m}(x)+R_{\lambda,m,\ell}(x),$ (3.61)
where we have used $\widehat{\chi_{m,j_{\ell}}}(\lambda^{\ell}a_{m})=1$. The remainder term $R_{\lambda,m,\ell}$ is bounded as
${\left\|R_{\lambda,m,\ell}\right\|}_{L^{p}}$ $\leqslant|b_{\ell}|{\left\|\nabla f_{m}\right\|}_{L^{p}}\int_{\mathbb{R}^{n}}|y||\chi_{m,j_{\ell}}(y)|dy$ (3.62)
$\leqslant C|b_{\ell}|2^{-j_{\ell}}{\left\|\nabla f_{m}\right\|}_{L^{p}}$ (3.63)
$=C|b_{\ell}|\lambda^{-\ell}{\left\|\nabla f_{m}\right\|}_{L^{p}}.$ (3.64)
@@
${\left\|P_{m}\Delta_{j_{\ell}}g_{\lambda,m,\ell}\right\|}_{L^{p}}$ $\geqslant|b_{\ell}|{\left\|f_{m}\right\|}_{L^{p}}-C\lambda^{-\ell}|b_{\ell}|{\left\|\nabla f_{m}\right\|}_{L^{p}}.$ (3.65)
By (3.57) and (3.65), it holds
${\left\|F_{\lambda,m,L}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{s}}\geqslant{}$ $c\left\{\sum_{\ell=1}^{L}\left(\lambda^{s\ell}|b_{\ell}|\right)^{q}\right\}^{\frac{1}{q}}{\left\|f_{m}\right\|}_{L^{p}}-C\sum_{\ell=1}^{L}\lambda^{(s-1)\ell}|b_{\ell}|{\left\|\nabla f_{k}\right\|}_{L^{p}}.$ (3.66)
On the estimate of $F^{({\rm r})}_{\lambda,m,L}$, the similar strategy as for the estimate of $F^{({\rm r})}_{\lambda,m,L}$ yields
${\left\|F^{({\rm r})}_{\lambda,m,L}\right\|}_{\dot{B}_{p,1}^{s}}\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j:(m,j)\in R_{k,\ell}}\sum_{|j^{\prime}-j|\leqslant 1}2^{sj^{\prime}}{\left\|P_{m}\Delta_{j^{\prime}}\Delta_{j}g_{\lambda,k,\ell}\right\|}_{L^{p}}$ (3.67)
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\sum_{j:(m,j)\in R_{k,\ell}}2^{sj}{\left\|P_{m}\Delta_{j}g_{\lambda,k,\ell}\right\|}_{L^{p}}$ (3.68)
$\leqslant{}$ $C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}}$ (3.69)
$+C\sum_{k=1}^{K}\sum_{\ell=1}^{L}\lambda^{(s+n(1-\frac{1}{p})-M)\ell}|b_{\ell}|{\left\|\nabla^{M}f_{k}\right\|}_{L^{p}\cap L^{1}}.$ (3.70)
Gathering (3.66) and (3.70), we obtain (3.9) and complete the proof. ∎